ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 120, 2000

Adolphson, A. and S. Sperber Exponential sums on A^{n}  3

Besser, A. Syntomic regulators and p-adic integration I: Rigid syntomic regulators

291

Besser, A. Syntomic regulators and p-adic integration II: K_{2} of curves

335

Canetti, R., J. Friedlander, S. Konyagin, M. Larsen, D. Lieman and I. Shparlinski On the statistical properties of Diffie– Hellman distributions

23

de Jong, A. J. and N. M. Katz Monodromy and the Tate Conjecture: Picard numbers and Mordell–Weil ranks in families

47

de Shalit, E. Foreword  289

Fouvry, É. Consequences of a result of N. Katz and G. Laumon concerning trigonometric sums

81

Friedlander, J. See R. Canetti

Heath-Brown, D. R. Kummer’s conjecture for cubic Gauss sums  97
Hida, H. Adjoint Selmer groups as Iwasawa modules  361

Huxley, M. N. and N. Watt The number of ideals in a quadratic field, II

125

Iovita, A. Formal sections and de Rham cohomology of semistable Abelian varieties

429

Iwaniec, H. and P. Sarnak The non-vanishing of central values of automorphic L-functions and Landau–Siegel zeros

155

Katz, N. M. See A. J. de Jong

Kim, H. H. and F. Shahidi Holomorphy of Rankin triple L-functions; special values and root numbers for symmetric cube L-functions

449

Konyagin, S. See R. Canetti

Kowalski, E. and P. Michel Explicit upper bound for the (analytic)

rank of J_{0}(q) 179

Kurlberg, P. The distribution of spacings between quadratic residues, II

205

Larsen, M. See R. Canetti

Laumon, G. Exponential sums and ℓ-adic cohomology: A survey  225

Lieman, D. See R. Canetti

Livné, R. Foreword  1

Michel, P. See E. Kowalski Ngô, B. C. Preuve d’une conjecture de Frenkel–Gaitsgory–Kazhdan–

Vilonen pour les groupes linéaires généraux

259

Panchishkin, A. A. On the Siegel–Eisenstein measure and its applications

467

Rudnick, Z. and A. Zaharescu The distribution of spacings between small powers of a primitive root

271

Sarnak, P. See H. Iwaniec Shahidi, F. See H. H. Kim Shparlinski, I. See R. Canetti

Soudry, D. Full multiplicativity of gamma factors for SO_{2ℓ+1} × GL_{n} 511

Sperber, S. See A. Adolphson Watt, N. See M. N. Huxley Zaharescu, A. See Z. Rudnick